4. Write as a product of a whole number and the logarithm
a) log(9)
b) log(25)
c) log(8)
d) log(27)
e) log(100)
f) log(32)
g) log(343)
h) log(128)
Can you please help me out? Thanks so much in advance:)
Can you please help me out? Thanks so much in advance:)

Answer by lwsshak3(11628)

4. Write as a product of a whole number and the logarithm
a) log(9)=log(3^2)=2log(3)
b) log(25)=log(5^2)=2log(5)
c) log(8)=1og(2^3)=3log(2)
d) log(27)=1og(3^3)=3log(3)
e) log(100)=1og(10^2)=2log(10)
f) log(32)=1og(2^5)=5log(2)
g) log(343)=1og(7^3)=3log(7)
h) log(128)=1og(2^7)=7log(2)